5 litre of a gas at a pressure of 2 atmospheric pressure and a temperature of 27 degree Celsius is heated to 227degree celcius and the pressure is raised to 6 atmospheric pressure . What is the volume under these conditions

1
Expert's answer
2022-01-24T13:39:05-0500

pV = nRT

R = 0.082 L atm / K x mol

T = 27 + 273 = 300 K

n = pV / RT = 2 x 5 / 0.082 x 300 = 0.4 mol

V1 = nRT / p = 0.4 x 0.082 x (227 + 273) / 6 = 2.7 L